# Bourgain’s Condition, Sticky Kakeya, and new Examples

### v1

@@ Abstract.
Our result follows from a new geometric characterization of Bourgain’s condition based on the structure of curved $\delta$-tubes in a $\delta^{1/2}$-tube. We give examples which show this property does not persist in a larger tube, and in particular in each dimension at least 3 there are operators satisfying Bourgain’s condition for which there is no diffeomorphism taking the corresponding family of curves to lines. This suggests that a general to sticky reduction in the spirit of Wang–Zahl needs substantial new ideas. We expect these examples to provide a good starting point.
@@ 1. Introduction
The operators $T_{\phi}^{\lambda}$ satisfy (1.4) if and only if $\phi$ satisfies Bourgain’s condition.
@@ 1.3. Sticky $\phi$-Kakeya sets, and a reduction to classical sticky Kakeya sets under Bourgain’s condition
It seems plausible that the sticky case is also the enemy case for Conjecture 1.6. The definition of stickiness for direction separated lines appearing in [wangZahlstickyKakeya] can be more naturally extended to $\phi$-Kakeya sets. To measure how closely $\phi$-curves pack together, we equip $\mathcal{C}(\phi)$ with the metric
@@
Theorem 1.17.
@@
In [nadjimzadah2025newcurvedkakeyaestimates] we announced, but did not prove anything about, a phase function in $\mathbb{R}^{3}$ satisfying Bourgain’s condition for which there does not exist a diffeomorphism taking the family of $\phi$-curves to lines. Below is a mild generalization of that example to all $n\geqslant 3$.
@@
We will prove Proposition 1.21 by showing that the union of curves through $\ell_{0,0}$ and a point $p\notin\ell_{0,0}$ is not contained in a surface, while of course a union of lines passing through a fixed line and point is contained in a plane. In [nadjimzadah2025newcurvedkakeyaestimates] this type of behavior is known as coniness, where it was exploited to give positive results for Kakeya sets of curves in a slightly different context. Figure 1.2 is a cartoon of this failure of (1.31) for $\phi_{n,\tan}$.
@@
Proposition 1.21 shows that the error in Theorem 1.8 cannot be strengthened to $O(|(\xi,v)-(\xi_{0},v_{0})|^{4})$, and we suspect that a version with error $O(|(\xi,v)-(\xi_{0},v_{0})|^{3})$ is also impossible. One should contrast Example 1.20 with translation-invariant phase functions and phase functions arising from Carleson–Sjölin operators on manifolds. If the phase function $\phi$ in either of those cases satisfies Bourgain’s condition, then $\mathcal{C}(\phi)$ is diffeomorphic to a family of lines [DaiOscillatory, gao2025curvedkakeyasetsnikodym]. Not only is $\mathcal{C}(\phi_{n,\tan})$ not diffeomorphic to a family of lines, which we believe is the first example satisfying Bourgain’s condition with this property, it is not even diffemorphic to a family of lines between scales $\delta$ and $\delta^{1/4}$.
Numerous arguments in Kakeya ultimately use the property that lines may be efficiently organized into planes. For example, Wolff’s hairbrush argument and the breakthrough of Wang–Zahl implicitely use this type of property. The curves in $\mathcal{C}(\phi_{n,\tan})$ do not organize into surfaces as lines organize into planes, so to go beyond the sticky case of Conjecture 1.6, new ideas are likely needed.
@@ 1.5. Organization of article
1.6. Potential and future work
Having proved the sticky case of Conjecture 1.6 in dimension 3, a natural next step is to attempt a general to sticky reduction in dimension 3 in the spirit of Wang–Zahl. Example 1.20 demonstrates that this will need some new ideas, such as perhaps polynomial Wolff axioms instead of convex Wolff axioms. We plan on attempting this approach, beginning with the $\tan$-example.
By following the proof of Theorem 1.19, the study of sticky $\phi$-Kakeya sets can be reduced to the study of certain semi-linearized sticky Kakeya sets, even when $\phi$ fails Bourgain’s condition. The curves in these semi-linearized sticky Kakeya sets take the form
@@ 1.6. Potential and future work
where $A(t)=tI_{n-1}+O(t^{2})$. Thus the study of sticky $\phi$-Kakeya sets seems significantly more tractable than general $\phi$-Kakeya sets. Theorem 1.8 shows that the structure of the semi-linearized families can completely describe Bourgain’s condition. It would be interesting to explore whether the semi-linearized families for other phase functions might suggest a more refined classification for Hörmander’s oscillatory integral problem. Some work in this direction based instead on the polynomial method and contact order appears in [DaiOscillatory].
@@ 1.7. Thanks
$(G(\mathbf{x},\xi_{0})\cdot\nabla_{\mathbf{x}})\nabla_{\xi}\phi(\mathbf{x},\xi_{0})=0$ (2.1)
@@ 6.2. The family of $\phi_{n,\tan}$-curves
We will obtain a contradiction as follows. We consider the $1$-parameter family of curves $\ell_{\xi(s),v(s)}$ passing through $\ell_{0,0}$ and a point $p\notin\ell_{0,0}$. We will show that the tangent directions $\gamma(s)$ of the curves $\ell_{s}$ at $p$ are not all contained in a 2-dimensional vector space. Concretely, we will show that $\gamma(s)\wedge\dot{\gamma}(s)\wedge\ddot{\gamma}(s)\neq 0$ for a fixed $s$. This is impossible by (6.20), since the lines passing through a fixed line and a point not on that line must be contained in a plane. To make the parameterization of the curves $\ell_{\xi(s),v(s)}$ easier to compute, we will first simplify the curves $\ell_{\xi,v}$ via a Taylor expansion and a diffeomorphism. Then we will write down the parameterization, and finally show that the tangent directions through $p$ are not contained in a $2$-plane. The details are as follows.
@@
We consider $s$ near 1 and solve for $\xi_{\mathbf{p}}(s)$ and $v_{\mathbf{p}}(s)$ to third order in $p$ near 0 (with $t_{0}>1$ fixed). We have $\xi_{0}(s)=v_{0}(s)=0$, so $\xi_{\mathbf{p}}(s),v_{\mathbf{p}}(s)=O(|p|)$. Below we will suppress the $\mathbf{p}$” in $\xi_{\mathbf{p}}(s)$ and $v_{\mathbf{p}}(s)$ for ease of notation. In coordinates, (6.25) gives the following system of equations
@@
Define $\gamma(s)=\frac{d}{dt}\ell_{\xi(s),v(s)}(t_{0})$, the tangent vector at point $p$ of the unique curve passing through $\ell_{0,0}(s)$ and $p$. We compute
@@
So that we don’t need to compute a tedious Gram determinant, choose $p^{\prime}=(0,\ldots,0,p_{n-2})$ so that $\gamma(1),\dot{\gamma}(1),\ddot{\gamma}(1)\in\{0\}^{n-3}\times\mathbb{R}^{3}$, and we may view these vectors as vectors in $\mathbb{R}^{3}$. We may finally compute
@@
By (6.20), $F(\tilde{\ell}_{\xi,v})$ is a family of lines. Thus for $s$ near 0, $F(\tilde{\ell}_{\xi(s),v(s)})$ is a family of lines passing through $F(\tilde{\ell}_{0,0})$ and $F(\mathbf{p})$, so it is contained in a fixed plane. In particular the tangent vectors of $F(\tilde{\ell}_{\xi(s),v(s)})$ at $F(\mathbf{p})$ lie in a fixed plane, so $\eta(s):=\frac{d}{dt}F(\ell_{\xi(s),v(s)}(t_{0}))$ satisfies $|\eta(1)\wedge\dot{\eta}(1)\wedge\ddot{\eta}(1)|=0$. But by the chain rule

### v2

@@ Abstract.
Our result follows from a new geometric characterization of Bourgain’s condition based on the structure of curved $\delta$-tubes in a $\delta^{1/2}$-tube. We find examples in all dimensions at least 3 which show this property does not persist in a larger tube, and in particular these are the first operators satisfying Bourgain’s condition for which there is no diffeomorphism taking the corresponding families of curves to lines. This suggests that a general to sticky reduction in the spirit of Wang–Zahl needs substantial new ideas. We expect these examples to provide a good starting point.
@@ 1. Introduction
The operators $T_{\phi}^{\lambda}$ satisfy (1.4) for $p>\frac{2n}{n-1}$ if and only if $\phi$ satisfies Bourgain’s condition.
@@ 1.3. Sticky $\phi$-Kakeya sets, and a reduction to classical sticky Kakeya sets under Bourgain’s condition
It seems plausible that the sticky case is also the enemy case for Conjecture 1.6. We choose to formalize stickiness for $\phi$-Kakeya sets by generalizing the definition in [wangZahlstickyKakeya], since it is more straightforward to write in general dimension. At least in $\mathbb{R}^{3}$, a suitable polynomial Wolff axiom version, analogous to the convex Wolff axiom version appearing in [wangZahlKakeya] can be adapted, along with a corresponding version of Theorem 1.19 below. We very briefly discuss this in Section 1.6. To measure how closely $\phi$-curves pack together, we equip $\mathcal{C}(\phi)$ with the metric
@@
Theorem 1.17 (Discretized classical sticky Kakeya in $\mathbb{R}^{3}$ [wangZahlstickyKakeya]).
@@
In each dimension $n\geqslant 3$ we give a phase function $\phi$ satisfying Bourgain’s condition, for which there does not exist a diffeomorphism taking the family of $\phi$-curves to lines. This answers a question of Gao–Liu–Xi [gao2025curvedkakeyasetsnikodym] in the negative, and we first announced (but did not prove anything about) this example in the case $n=3$ in [nadjimzadah2025newcurvedkakeyaestimates].
@@
We will prove Proposition 1.21 by showing that the union of curves through $\ell_{0,0}$ and a point $\mathbf{p}\notin\ell_{0,0}$ is not contained in a surface, while of course a union of lines passing through a fixed line and point is contained in a plane. In [nadjimzadah2025newcurvedkakeyaestimates] this type of behavior is known as coniness, where it was exploited to give positive results for Kakeya sets of curves in a slightly different context. Figure 1.2 is a cartoon of this failure of (1.31) for $\phi_{n,\tan}$.
@@
Proposition 1.21 shows that the error in Theorem 1.8 cannot be strengthened to $O(|(\xi,v)-(\xi_{0},v_{0})|^{4})$, and we suspect that a version with error $O(|(\xi,v)-(\xi_{0},v_{0})|^{3})$ is also impossible. One should contrast Example 1.20 with translation-invariant phase functions and phase functions arising from Carleson–Sjölin operators on manifolds. If the phase function $\phi$ in either of those cases satisfies Bourgain’s condition, then $\mathcal{C}(\phi)$ is diffeomorphic to a family of lines [DaiOscillatory, gao2025curvedkakeyasetsnikodym]. Not only is $\mathcal{C}(\phi_{n,\tan})$ not diffeomorphic to a family of lines, which we believe is the first example satisfying Bourgain’s condition with this property, it is not even diffeomorphic to a family of lines between scales $\delta$ and $\delta^{1/4}$.
Numerous arguments in the study of Kakeya sets ultimately use the property that lines may be efficiently organized into planes. For example, Wolff’s hairbrush argument uses that the $\delta$-tubes through a fixed stem can be organized into essentially disjoint slabs [wolffHairbrush]. The argument of Wang–Zahl uses that $\delta$-tubes can be organized into planks [wangZahlKakeya]. At least outside a fixed $(\delta^{1/4},\phi_{n,\tan})$-tube, the $(\delta,\phi_{n,\tan})$-tubes do not have properties analogous to these. Thus to go beyond the sticky case of Conjecture 1.6, new ideas are needed.
@@ 1.5. Organization of article
1.6. Discussion and future work
Having proved the sticky case of Conjecture 1.6 in dimension 3, a natural next step is to attempt a general to sticky reduction in dimension 3 in the spirit of Wang–Zahl. As mentioned above, in light of the $\tan$-example, this will require new ideas. The most promising route seems to be to replace the Katz-Tao and Frostman non-concentration conditions in convex sets [wangZahlKakeya], with the analogous non-concentration conditions in semi-algebraic sets with bounded complexity. Following the outline in [guth2025outlinewangzahlproofkakeya], a collection of $(\delta,\phi)$-tubes is then sticky if it satisfies the Katz–Tao polynomial Wolff axioms along a sequence of nearby scales. The version of Conjecture 1.13 in dimension $3$ with this notion of stickiness continues to be true by running an induction on scales argument similar to the proof of Theorem 1.19. Though one must adjust the inductive hypothesis accordingly and replace the application of Theorem 1.17 from [wangZahlstickyKakeya] with [wangZahlKakeya, Theorem 10.2]. The general to sticky reduction for Conjecture 1.6 will then require a better understanding of the grains that can arise from $(\delta,\phi)$-tubes. We are currently trying to understand these new phenomena for the $\tan$-example.
By following the proof of Theorem 1.19, even when $\phi$ is a phase function failing Bourgain’s condition, the study of sticky $\phi$-Kakeya sets can be reduced to the study of sticky $(\delta,\phi)$-tubes in a $(\delta^{1/2},\phi)$-tube. By Taylor expanding, we need to understand sticky Kakeya sets of certain semi-linearized families of curves of the form
@@ 1.6. Discussion and future work
where $A(t)=tI_{n-1}+O(t^{2})$. Theorem 1.8 shows that the structure of the semi-linearized families can completely describe Bourgain’s condition. It would be interesting to explore whether the semi-linearized families for other phase functions might suggest a more refined classification for Hörmander’s oscillatory integral problem. Some work in this direction based instead on the polynomial method and contact order appears in [DaiOscillatory].
@@ 1.7. Thanks
$(G(\mathbf{x},\xi)\cdot\nabla_{\mathbf{x}})\nabla_{\xi}\phi(\mathbf{x},\xi)=0$ (2.1)
@@ 6.2. The family of $\phi_{n,\tan}$-curves
We will obtain a contradiction as follows. We consider the $1$-parameter family of curves $\tilde{\ell}_{\xi(s),v(s)}$ passing through $\tilde{\ell}_{0,0}$ and a point $\mathbf{p}\notin\tilde{\ell}_{0,0}$. We will show that the tangent directions $\gamma(s)$ of the curves $\tilde{\ell}_{s}$ at $\mathbf{p}$ are not all contained in a 2-dimensional vector space. Concretely, we will show that $\gamma(s)\wedge\dot{\gamma}(s)\wedge\ddot{\gamma}(s)\neq 0$ for a fixed $s$. This is impossible by (6.20), since the lines passing through a fixed line and a point not on that line must be contained in a plane. To make the parameterization of the curves $\tilde{\ell}_{\xi(s),v(s)}$ easier to compute, we will first simplify the curves $\tilde{\ell}_{\xi,v}$ via a Taylor expansion and a diffeomorphism. Then we will write down the parameterization, and finally show that the tangent directions through $\mathbf{p}$ are not contained in a $2$-plane. The details are as follows.
@@
We consider $s$ near 1 and solve for $\xi_{\mathbf{p}}(s)$ and $v_{\mathbf{p}}(s)$ to third order in $p$ near 0 (with $t_{0}>1$ fixed). We have $\xi_{0}(s)=v_{0}(s)=0$, so $\xi_{\mathbf{p}}(s),v_{\mathbf{p}}(s)=O(|p|)$. Below we will suppress the $\mathbf{p}$” in $\xi_{\mathbf{p}}(s)$ and $v_{\mathbf{p}}(s)$ for ease of notation. In coordinates, (6.25) and (6.26) give the following system of equations
@@
Define $\gamma(s)=\frac{d}{dt}\tilde{\ell}_{\xi(s),v(s)}(t_{0})$, the tangent vector at point $\mathbf{p}$ of the unique curve passing through $\tilde{\ell}_{0,0}(s)$ and $\mathbf{p}$. We compute
@@
Choose $p^{\prime}=(0,\ldots,0,p_{n-2})$ so that $\gamma(1),\dot{\gamma}(1),\ddot{\gamma}(1)\in\{0\}^{n-3}\times\mathbb{R}^{3}$, and we may view these as vectors in $\mathbb{R}^{3}$. We may finally compute
@@
By (6.20), $F(\tilde{\ell}_{\xi,v})$ is a family of lines. Thus for $s$ near 0, $F(\tilde{\ell}_{\xi(s),v(s)})$ is a family of lines passing through $F(\tilde{\ell}_{0,0})$ and $F(\mathbf{p})$, so it is contained in a fixed plane. In particular the tangent vectors of $F(\tilde{\ell}_{\xi(s),v(s)})$ at $F(\mathbf{p})$ lie in a fixed plane, so $\eta(s):=\frac{d}{dt}F(\tilde{\ell}_{\xi(s),v(s)}(t_{0}))$ satisfies $|\eta(1)\wedge\dot{\eta}(1)\wedge\ddot{\eta}(1)|=0$. But by the chain rule